OpenStudy (anonymous):

Find the mean, median, and mode of the data set. Round to the nearest tenth. test scores on a math exam: 83, 96, 67, 67, 86, 65, 82, 85, 97, 73, 85, 67, 99, 98, 65, 96, 70, 92, 74, 83, 72, 76 a. mean = 74.1, median = 82.5, mode = 67 b. mean = 74.1, median = 82.5, mode = 82 c. mean = 80.8, median = 82.5, mode = 67 d. mean = 80.8, median = 82, mode = 67

OpenStudy (texaschic101):

first...put your numbers in order.. the mean is the average....add up all the numbers then divide by how many numbers there are. the median is the middle number the mode is the number that appears the most. 65, 65, 67 , 67 , 67 , 70 , 72 , 73 , 74, 76 , 82, 83, 83, 85, 85, 86, 92, 96, 96, 97, 98, 99 mean = 1778/22 = 80.8 median = (82 + 83) / 2 = 165/2 = 82.5 mode = 67 questions ?

OpenStudy (anonymous):

no and thank you